# Phonon-mediated electron attraction in SrTiO3 via the generalized Fröhlich and deformation potential mechanisms

### v1

@@ Abstract
To date, the precise mechanism of the pairing between electrons in $\text{SrTiO}_{3}$ remains elusive. As $\text{SrTiO}_{3}$ is a material with strong electron-phonon coupling [12, 13], several works have investigated the possibility of phonon-mediated superconductivity [1, 14, 15, 16, 17, 18, 19]. Proposed mechanisms include inter-valley scattering [1, 14], coupling to polar [15], and defect [16] phonons, higher-order electron-phonon interactions [17], coupling to the soft infrared active zone-center phonon of $\text{SrTiO}_{3}$ [18], Rashba coupling [19], and more. However, as also summarized in recent reviews [20, 21], the emergence of superconductivity through these mechanisms generally depends on adjustable parameters that are challenging to compute from first principles or to accurately infer from experiment. A fully first-principles explanation for the phonon-mediated electron-electron attraction has not been possible to achieve, as $\text{SrTiO}_{3}$ becomes superconducting at low doping densities where the Fermi energy is much smaller than the phonon frequencies, placing it in the non-adiabatic limit. As a result, the widely used Migdal-Eliashberg framework for phonon-mediated superconductivity does not apply [22], leading to an ongoing debate on the mechanism of Cooper pair formation.
Here we present a theoretical framework to rigorously incorporate lattice screening in electron-electron interactions. Without relying on adjustable parameters, our theory analytically demonstrates that linear electron-phonon coupling to multiple longitudinal optical (LO) phonons, alone is sufficient to yield long-range attraction between electrons in $\text{SrTiO}_{3}$. Moreover, we translate our theory to an efficient computational scheme, and show that the deformation potential associated with vibrations of the mixed ionic-covalent Ti-O bond provides an additional strong, short-range attractive contribution to electron-electron interactions. Our theory, outlined in detail in our companion paper [23], applies to several doped semiconductors and insulators, and our predicted characteristics of the phonon-mediated electron attraction in $\text{SrTiO}_{3}$ are consistent with the qualitative features of its superconductivity.
@@
where we have suppressed the band index for simplicity. Here we neglect the screening of the electronic Coulomb interaction by free carriers in doped systems, which is justified given the low doping densities at which $\text{SrTiO}_{3}$ becomes superconducting. For the phonon-mediated interaction, as detailed in our companion paper [23], we obtain for periodic systems
@@
In polar semiconductors and insulators, dipolar electron-phonon coupling is the dominant term in the electron-phonon interaction [30]. This long-range electron-phonon coupling is written as
@@
This expression was used by Gor’kov [15] in an effort to explain the emergence of attractive electron-electron interactions in $\text{SrTiO}_{3}$. However, even in the extreme case of $\epsilon_{0}\rightarrow\infty$, this will at best cancel out the repulsive electron-electron Coulomb interaction of eq. (3), and cannot result in a net attraction [33]. This later led Gor’kov to hypothesize a mechanism of electrons coupling to defect phonons in order to provide the missing additional attraction [16]. However, we see that our generalized expression of eq. (7), that does not rely on the aforementioned simplifications, provides a distinct answer: long-range electron-phonon coupling alone is sufficient to yield a strong attraction, once the effect of multiple LO phonons is accounted for fully.
@@
Acknowledgments.–The authors are thankful to Marvin L. Cohen for helpful and inspiring discussions. This material is based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems Center (SQMS) under contract No. DE-AC02-07CH11359. We are grateful for support from NASA Ames Research Center. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC awards HEP-ERCAP0029167 and DDR-ERCAP0029710.

### v2

@@ Abstract
To date, the precise mechanism of the pairing between electrons in $\text{SrTiO}_{3}$ remains elusive. As $\text{SrTiO}_{3}$ exhibits strong electron-phonon coupling [12, 13], several works have investigated the possibility of phonon-mediated superconductivity [1, 14, 15, 16, 17, 18, 19]. Proposed mechanisms include inter-valley scattering [1, 14], coupling to polar [15], and defect [16] phonons, higher-order electron-phonon interactions [17], coupling to the soft infrared active zone-center phonon of $\text{SrTiO}_{3}$ [18], which is enhanced by anharmonic damping [20], and Rashba coupling [19, 21]. However, as also summarized in recent reviews [22, 23], the emergence of superconductivity through these mechanisms generally depends on adjustable parameters that are challenging to compute from first principles or to accurately infer from experiment. A fully first-principles explanation for the phonon-mediated electron-electron attraction remains elusive, as $\text{SrTiO}_{3}$ becomes superconducting at low doping densities where the Fermi energy is much smaller than the phonon frequencies, placing it in the non-adiabatic limit. As a result, the widely used Migdal-Eliashberg framework for phonon-mediated superconductivity does not apply [24], leading to an ongoing debate on the mechanism of Cooper pair formation.
Here we present a theory that rigorously incorporates lattice screening in electron-electron interactions. Without relying on adjustable parameters, we analytically demonstrate that linear electron-phonon coupling to multiple longitudinal optical (LO) phonons, alone is sufficient to yield long-range attraction between electrons in $\text{SrTiO}_{3}$. Moreover, we translate our theory to a first-principles computational scheme, and show that the deformation potential associated with vibrations of the Ti-O bond provides an additional strong, short-range attractive contribution to electron-electron interactions. Our theory, outlined in detail in our companion paper [25], applies to several doped semiconductors and insulators, and our predicted characteristics of the phonon-mediated electron attraction in $\text{SrTiO}_{3}$ are consistent with the qualitative features of its superconductivity.
@@
where we have suppressed the band index, and neglected the screening of the Coulomb interaction by free carriers in doped systems, which is justified given the low doping densities at which $\text{SrTiO}_{3}$ becomes superconducting. For the phonon-mediated interaction, as detailed in our companion paper [25], we obtain for periodic systems
@@
In polar semiconductors and insulators, long-range dipolar electron-phonon coupling dominates the electron-phonon interaction [32], and is written as
@@
This expression was used by Gor’kov [15] in an effort to explain the emergence of attractive electron-electron interactions in $\text{SrTiO}_{3}$. However, even in the extreme case of $\epsilon_{0}\rightarrow\infty$, this will at best cancel out the repulsive electron-electron Coulomb interaction of eq. (3), and cannot result in a net attraction [35]. This later led Gor’kov to hypothesize that electron coupling to defect phonons provides the missing additional attraction [16]. However, we see that our generalized expression of eq. (7), that does not rely on the aforementioned simplifications, provides a distinct answer: long-range electron-phonon coupling alone is sufficient to yield a strong attraction, once the effect of multiple LO phonons is accounted for fully.
@@
Acknowledgments.–The authors are thankful to Marvin L. Cohen for helpful and inspiring discussions. This material is based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems Center (SQMS) under contract No. DE-AC02-07CH11359. We are grateful for support from NASA Ames Research Center. VV was supported by the National Science Foundation (NSF) CAREER award through grant No. DMR-1945098. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC awards HEP-ERCAP0029167 and DDR-ERCAP0029710.